survey of magazine subscribers showed that 45.8% rented a car during the past 12 months for business reasons, 54% rented a car during the past

12 months for personal reasons, and 30% rented a car during the past 12 months for both business and personal reasons.

i. What is the probability that a subscriber rented a car during the past 12 months for business or personal reasons?

ii. What is the probability that a subscriber did not rent a car during the past 12 months for either business or personal reasons




1
Expert's answer
2021-09-14T04:24:01-0400

QUESTION

survey of magazine subscribers showed that 45.8% rented a car during the past 12 months for business reasons, 54% rented a car during the past 12 months for personal reasons, and 30% rented a car during the past 12 months for both business and personal reasons.

i. What is the probability that a subscriber rented a car during the past 12 months for business or personal reasons?

ii. What is the probability that a subscriber did not rent a car during the past 12 months for either business or personal reasons?

SOLUTION

We can solve this problem by building the "Venn Diagram" of these probabilities.

I am going to assume that A is the probability that a magazine subscriber rented a car for business reasons and B is the probability that a magazine subscriber rented a car for personal reasons and C are those who did not rent a car for either of these reasons.

So we would have: "A=a + (A\bigcap B)"

Where:

"a" represent those who only rented for business reasons.

"(A\bigcap B)" represent those who rented both for business and personal reasons.

From the same logic, we have that: "B = b + (A \bigcap B)", Where:

"b" represent those who only rented for personal reasons.

The sum of the probabilities is 1, so:

"a+b+(A\bigcap B)+C=1"

Begin getting the values from the intersection of these sets.

30% of the people rented a car during the past 12 months for both business and personal reasons. So, "(A\bigcap B)=0.3"

45.8% rented a car during the past 12 months for business reasons, meaning that "A=0.458"

and "A=a+(A\bigcap B)"

"0.458=a+0.3"

"a=0.158"

54% rented a car during the past 12 months for personal reasons.

"B=b+(A\bigcap B)" Meaning that "B=0.54"

And "B=b+(A\bigcap B)"

"0.54=b+0.3"

"b=0.21"

SO WE MOVE TO SOLVING THE QUESTIONS:

i. What is the probability that a subscriber rented a car during the past 12 months for business or personal reasons?

Solution:

That is the probability that someone rented a car for only one of these reasons, or both.

So: "P=a+b+(A\bigcap B)"

"P=0.158+0.24+0.3"

"P=0.698"

69.8% probability that a subscriber rented a car during the past 12 months for business or personal reasons.

ANSWER: "69.8\%"

ii. What is the probability that a subscriber did not rent a car during the past 12 months for either business or personal reasons?

Solution:

This is the value of C, so we have that: "a+b+(A\bigcap B)+C=1"

"0.698+C=1"

"C=0.302"

30.2% probability that a subscriber did not rent a car during the past 12 months for either business or personal reasons.

ANSWER: "30.2\%"